OpenStudy (anonymous):

Find an equation in standard form of the line through p(-5,2) and having the slope -3

OpenStudy (zepp):

Standard like ax+b?

OpenStudy (zepp):

Umm, sorry, was away :( The formula to find a straight-line equation is y – y1 = m(x – x1)

OpenStudy (zepp):

Where m is the slope and x1 and y1 is the point. (x,y)

OpenStudy (zepp):

So you do y - (2) = (-3)(x - (-5)) y - 2 = -3 (x+5) y - 2 = -3x + -15 y - 2 = -3x -15 y = -3x - 15 + 2 y = -3x - 13 :))

OpenStudy (zepp):

And it's important to always verify your equation every time, you just have to plug the x from the point and find out if y is the same as the y from the point. y = -3(-5) -13 y = 15 -13 y = 2 So the equation is right.